OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 4x + 2y = 5 x – 3y = 7 –14 –29 23 –1

OpenStudy (amistre64):

-(5) 45 17 +(28)

OpenStudy (amistre64):

wait ... thats the y i think

OpenStudy (amistre64):

now i gotta google the terminology to remember which D is which lol

OpenStudy (amistre64):

Dx uses y parts .... ugh

OpenStudy (amistre64):

-(-15) 2 5 -3 7 +(14) im gonna assume its 29 unless told otherwise :)

OpenStudy (amistre64):

fine .. -29

OpenStudy (anonymous):

thank you so much for your help :)

OpenStudy (anonymous):

\[\left|\left( \begin{array}{cc} 5 & 2 \\ 7 & -3 \\ \end{array} \right)\right|=-29\]

OpenStudy (anonymous):

If you were asked to find x \[x==\frac{\left|\left( \begin{array}{cc} 5 & 2 \\ 7 & -3 \\ \end{array} \right)\right|} {\left|\left( \begin{array}{cc} 4 & 2 \\ 1 & -3 \\ \end{array} \right)\right|}=\frac{-29}{-14}=\frac{29}{14}\]